OpenStudy (anonymous):

What expression is equivalent to? (Points : 1) a ≥ –4 or a ≤ 4 a ≤ –4 or a ≥ 4 a ≥ –4 and a ≤ 4 a ≤ –4 and a ≥ 4

OpenStudy (radar):

I can see 4 expressions, but I don't see the equivalent expression. Is all of the problem there?

OpenStudy (anonymous):

|a| < 4 but there's the equal line under it

OpenStudy (radar):

O.K. The absolute value of a is less than 4 Hmmm let think about this.

OpenStudy (radar):

It will be one of the "and" choices. Read them both and remember that for the AND both of the conditions must be met at the same time.

OpenStudy (anonymous):

|a| le 4

OpenStudy (radar):

Less than or equal\[a \le 4\]

OpenStudy (anonymous):

do I flip the sign?

OpenStudy (radar):

No, just look at the two "and" statements and see which one could work. As an example looking at the last one \[a \le -4 and a \ge 4\] Note that the absolute value of a could be 6 which does not hold true for the problem. However the third option a can only range from -4 to +4 which would satisfy the problem.

OpenStudy (radar):

*anda should be a space between d and a. The last option allows values for a that range from -4 to - infinity but the and them with values that range from positive 4 to infinity therefore it is an inconsistent statement.

OpenStudy (radar):

Good luck with your studies.